### v1

# Optical probes of two-component pairing states in transition metal dichalcogenides

Abstract
Certain transition metal dichalcogenide heterostructures based on 2H-NbSe2 and 4Hb-TaS2 have recently displayed surprising signatures of unconventional superconductivity. While the pairing channel remains unknown, it has been argued that spin fluctuations should lead to a two-component $E^{\prime}$ pairing state which is compatible with some experimental features. Exploiting the particular multi-orbital character of the Fermi surface and the presence of Ising spin-orbit coupling which enable finite optical conductivity in the clean limit, in this work we predict clear-cut optical signatures of the chiral and nematic ground states of the $E^{\prime}$ pairing. In particular, we show how the spontaneous breaking of time-reversal and threefold symmetry are reflected in the optical Hall and anisotropic conductivities, respectively, while different spectral features can be connected with the momentum dependence of the gap functions. Our work provides a fingerprint that can be measured experimentally, constraining the pairing channel in TMD superconductors, and will help determine whether superconductivity is topological in these systems.
Introduction–– Metallic transition metal dichalcogenides (TMDs) MX2 (M=Nb,Ta, X=S, Se) Wilson et al. (1975), believed to be standard BCS superconductors in their bulk 2H polytype, have recently shown signatures of unconventional superconductivity Norman (2011) when made into thin films and heterostructures, calling for a better understanding of their pairing symmetries and mechanisms. For example, NbSe2 thin films display spontaneous breaking of threefold symmetry in the superconducting state under in-plane magnetic fields (indicative of a two-component order parameter) Hamill et al. (2021); Cho et al. (2022), as well as superconducting collective modes Wan et al. (2022) and anomalous resilience to magnetic fields Kuzmanović et al. (2022). Topological superconductivity, necessarily unconventional, has also been reported in monolayer CrBr3 on bulk NbSe2 Kezilebieke et al. (2020). 4Hb-TaS2 Di Salvo et al. (1973), which can be seen as a heterostructure of H and T polytypes, shows spontaneous breaking of time-reversal symmetry (TRS) at the critical superconducting temperature $T_{c}$ Ribak et al. (2020), spontaneous vortices in the superconducting state Persky et al. (2022), superconducting edge modes Nayak et al. (2021) and transport evidence of a two-component order parameter, which also includes spontaneous breaking of threefold symmetry Almoalem et al. (2022); Silber et al. (2024). A chiral superconducting state has also been reported in 2H-TaS2 intercalated with chiral molecules Wan et al. (2023).
The spontaneous breaking of threefold or time-reversal symmetries suggest a two-component SC order parameter, which may condense into nematic or chiral states depending on the experiment, possibly competing with a conventional pairing channel. The microscopic origin of such two-component pairing remains unknown, but it is strongly constrained by the presence of Ising spin-orbit coupling (SOC) Xi et al. (2016); Sohn et al. (2018); De la Barrera et al. (2018); Wickramaratne et al. (2020), induced by a crystal structure which breaks inversion symmetry but preserves a horizontal mirror $\sigma_{h}$. The pairing problem for H-TMDs has been studied at length He et al. (2018); Möckli and Khodas (2018); Shaffer et al. (2020); Margalit et al. (2021); Dentelski et al. (2021); Hörhold et al. (2023); Liu et al. (2024), in particular in the presence of two-component order parameters Hsu et al. (2017); Chen et al. (2019); Lane and Zhu (2022), but without identifying a microscopic interaction that makes these channels attractive. Previous works had however identified that monolayer H-TMDs are near a magnetic instability towards an antiferromagnetic state with in-plane polarization Zheng et al. (2018); Wickramaratne et al. (2020); Das and Mazin (2021); Costa et al. (2022), which leads to strong transverse paramagnons that can mediate unconventional superconductivity. A more recent calculation reveals that such fluctuations favor an $E^{\prime}$ triplet state in the presence of Ising SOC Roy et al. (2024); Siegl et al. (2024), although it remains to be established whether this solution is robust in the presence of additional electron-phonon coupling Das et al. (2023). This $E^{\prime}$ channel, which pairs electrons of opposite spins (and is therefore unaffected by Ising SOC), is a two-component representation of the point group that can condense in nematic $(p_{x},p_{y})$ or chiral $p_{x}+ip_{y}$ ground states. It is thus a strong candidate to explain the phenomenology observed in TMD systems, and an experimental probe that can distinguish it from other channels would be desirable.
Figure 1: H-polytype of TMDs. (a) Crystalline structure of a single H-layer, corresponding to an ABA stacking of hexagonal lattices. (b) Band structure of the normal state of the H layer of TaS2. (c) Optical conductivity in the normal state of TaS2 shown in the range where it is nonzero.
It has recently been realized that the optical conductivity is a useful probe of the superconducting state of multiband (or multiorbital) systems Ahn and Nagaosa (2021a). While in single-band superconductors the optical conductivity is essentially disorder-driven and vanishes in the clean limit, clean multiband superconductors have finite optical conductivity, which is sensitive to the pairing channel and is connected to the superfluid weight Ahn and Nagaosa (2021b); Chen and Huang (2021). It is also ideally suited to probe the difference between nematic and chiral ground states. Conductivity anisotropy $\sigma_{xx}-\sigma_{yy}$ is sensitive to the breaking of $C_{3}$ symmetry Xu et al. (2022), while optical Hall conductivity is only allowed in chiral systems which break time-reversal symmetry Taylor and Kallin (2012); Wang et al. (2014, 2017). Previous work has considered the non-linear optical response of non-centrosymmetric superconductors Xu et al. (2019); Tanaka et al. (2023), and recently both linear and non-linear conductivities were computed for the H-TMD three band model Raj et al. (2024) with the single-component pairing states proposed in Ref. [Margalit et al., 2021]. The $E^{\prime}$ state, currently a strong candidate for unconventional pairing Roy et al. (2024); Siegl et al. (2024), was not considered.
In this work, we consider the optical conductivity of the $E^{\prime}$ superconducting state in both its chiral and nematic ground states, demonstrating that the optical Hall conductivity $\sigma_{xy}^{H}$ and the conductivity anisotropy $\sigma_{xx}-\sigma_{yy}$, which are zero for all other superconducting states, serve as unique probes for the chiral and nematic $E^{\prime}$ states, respectively.
Model and pairing channels–– The H polytype of TMDs MX2 (M=Nb,Ta, X=S, Se) crystallizes in a trigonal prismatic crystal structure (Fig 1 (a)) with $D_{3h}$ point-group symmetry. The Fermi surface derives from the ${d_{z^{2}},d_{x^{2}-y^{2}},d_{xy}}$ orbitals of the transition metalMöckli and Khodas (2018). We consider a spinful three-band tight-binding model with up to third-nearest neighbours hoppingsLiu et al. (2013) in the orbital basis $\{d_{z^{2}},d_{x^{2}-y^{2}},d_{xy}\}$ which reads
$h_{\bm{k}}=\left[\begin{array}[]{cc}h^{\mathrm{TNN}}_{\bm{k}}+\frac{\lambda}{2}L_{z}&0\\ 0&h^{\mathrm{TNN}}_{\bm{k}}-\frac{\lambda}{2}L_{z}\end{array}\right],$ (3)
where $h^{\mathrm{TNN}}_{\bm{k}}$ is the tight-binding Hamiltonian with up to third nearest-neighbours hoppings, $\lambda$ is the SOC strength, and $L_{z}$ is the matrix of the z-component of the angular momentum in the orbital basis $\{d_{z^{2}},d_{x^{2}-y^{2}},d_{xy}\}$ (See Ref. [Liu et al., 2013] for details and Fig. 1 (b) for the resulting band structure.).
Superconductivity is described with the Bogoliubov-de Gennes (BdG) Hamiltonian defined in terms of Nambu spinors $\Psi=(\psi,\psi^{\dagger})^{T}$ as
$\mathcal{H}=\frac{1}{2}\Psi^{\dagger}H_{\rm BdG}\Psi=\frac{1}{2}\Psi^{\dagger}_{\bm{k}}\left(\begin{array}[]{cc}h_{\bm{k}}&\Delta_{\bm{k}}\\ \Delta_{\bm{k}}^{\dagger}&-h_{-\bm{k}}^{*}\end{array}\right)\Psi_{\bm{k}}$ (6)
with $\Delta^{\dagger}_{\bm{k}}=-\Delta_{-\bm{k}}^{*}$ the superconducting gap. For simplicity we consider only momentum-independent pairing states where $\Delta_{\bm{k}}=\Delta$ is a 6x6 matrix in spin and orbital space.
The different pairing channels are classified as irreducible representations (irreps) of the $D_{3h}=M_{z}\otimes C_{3v}$ point group, generated by mirror planes perpendicular to the $x$-axis $M_{x}$ and $z$-axis $M_{z}$, and threefold rotations $C_{3}$ around the $z$-axis. This group contains the $M_{z}$-even irreps $A_{1}^{\prime}$, $A_{2}^{\prime}$, $E^{\prime}$ and $M_{z}$-odd irreps $A_{1}^{\prime\prime}$, $A_{2}^{\prime\prime}$, $E^{\prime\prime}$. Since the orbitals that contribute to the bands at the Fermi level are all $M_{z}$-even and we only consider in-plane pairing, odd $M_{z}$ parity can only come from the spin part of the pairing. Singlet pairing $\Delta=i\sigma_{y}$ and out-of-plane triplet $\Delta=i\sigma_{y}\sigma_{z}=\sigma_{x}$ are $M_{z}$-even pairings (which correspond to opposite spin pairing), while in-plane triplets $\Delta=i\sigma_{y}(\sigma_{x},\sigma_{y})$ are $M_{z}$-odd pairings (which correspond to equal-spin pairing). Since $D_{3h}$ does not contain the inversion operator, singlet and triplet channels can generally mix due to SOC, and channels are only classified by the global (spin and orbital) symmetry of the pairing.
Previous works have discussed different unconventional equal-spin pairing channels driven by the breaking of $M_{z}$. An equal-spin pairing chiral p+ip superconducting state was predicted in the presence of large Rashba spin-orbit coupling Shaffer et al. (2020), while Ref. [Margalit et al., 2021] proposed an equal-spin pairing one-dimensional irrep $A_{2}^{\prime\prime}$. Extending beyond these works, Ref. [Liu et al., 2024] proposed an interlayer spin-polarized pairing state with odd parity. Given the large Ising SOC, which suppresses equal-spin pairing, $M_{z}$-even states are however more likely to be realized unless the scale of $M_{z}$ breaking is as large as Ising SOC. Although the $A_{1}^{\prime}$ channel, which corresponds to the standard $s$-wave state (potentially with a symmetry-allowed triplet admixture), is the most common $M_{z}$-even state favored by electron-phonon interactions, recent works show that strong spin fluctuations in this system should favor the $E^{\prime}$ channel Roy et al. (2024); Siegl et al. (2024).
In addition to being unaffected by Ising SOC and favored by spin fluctuations, the $E^{\prime}$ channel has a number of appealing features. First, as a two-component order parameter it admits both chiral and nematic ground states, in line with the phenomenology in 4Hb-TaS2. In addition, this state has interesting topological features. $M_{z}$-even pairing with Ising SOC implies spin rotation invariance around $z$ direction, which puts this system in topological class AIII Schnyder et al. (2008), featuring $Z$ invariants in 1D that can protect nodal points in the case of nematic ground states. For chiral ground states, each mirror sector realizes a Chern number $C=3$ superconductor. In the case of triplet $E^{\prime}$ chiral state, a $C=6$ superconductor is realized, which could also be consistent with the edge state phenomenology in 4Hb-TaS2Nayak et al. (2021) and CrBr3/NbSe2Kezilebieke et al. (2020).
Figure 2: Band structure of the BdG Hamiltonian in Eq.(6) for 4Hb-TaS2 with the superconducting pairings (a) $\Delta_{p_{x}}$, (b) $\Delta_{p_{y}}$, and (c) $\Delta_{p+ip}$ for $\Delta=0.1$ eV. The dark (light) gray shadowing corresponds to the minimal (maximal) size of the gap in the band structure along the high-symmetry directions. The insets correspond to the gap size across the Brillouin Zone for each pairing, where lightest colors are closer to zero gap, and indicate the high-symmetry directions corresponding to the band structure plots. The dotted-dashed line in (c) corresponds to the energy scale associated with the Van Hove singularity giving rise to the secondary peak in $\sigma_{xx}$ and $\sigma^{H}_{xy}$.
The discussion above motivates us to consider the triplet $E^{\prime}$ channel, given in the orbital basis $\{d_{z^{2}},d_{xy},d_{x^{2}-y^{2}}\}$ as the two-component basis pairing matrices
$\{\Delta_{p_{x}},\Delta_{p_{y}}\}=\sigma_{y}\sigma_{z}\otimes\left\{{\left(\begin{array}[]{ccc}0&i&0\\ -i&0&0\\ 0&0&0\end{array}\right)},{\left(\begin{array}[]{ccc}0&0&-i\\ 0&0&0\\ i&0&0\end{array}\right)}\right\}.$ (7)
This term represents spin-triplet, orbital singlet pairing between the out-of-plane orbital $d_{z^{2}}$ and each of the in-plane orbitals $d_{x^{2}-y^{2}}$ and $d_{xy}$. The superconducting phases with pairings $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$ represent nematic phases, while the chiral superconducting phase is given by the pairing $\Delta_{p+ip}=\Delta_{p_{x}}+i\Delta_{p_{y}}$.
BdG quasiparticle band structure –– For concreteness we consider a single layer of H-TaS2 and use the normal state tight-binding parameters reported in Ref. [Margalit et al., 2021]. Since to resolve band structure and optical features with realistic values of the superconducting gap ($\sim$1 meV) would require a very fine k-mesh with high computational cost, for illustrative purposes we consider pairings of an artificially large strength $\Delta=0.1$ eV, keeping in mind that experimental features are to be expected at the real values of the superconducting gap. The three different ground states with $E^{\prime}$ pairing $\Delta_{p_{x},p_{y},p+ip}$ lead to the three different band structures in Fig. 2. The $\Delta_{p_{x}}$ pairing is gapless along the $\Gamma M^{\prime}$ ($k_{x}=0$) direction, while the $\Delta_{p_{y}}$ pairing is gapless along the $\Gamma K$ ($k_{y}=0$) direction. In both cases we see that $\Gamma M$ and $\Gamma M^{\prime}$ are not equivalent, a consequence of the breaking of $C_{3}$ symmetry. For both nematic phases, the maximum gap is approximately $1.6\Delta$ along the $\Gamma K$ ($\Gamma M^{\prime}$) directions for $\Delta_{p_{x}}$ ($\Delta_{p_{y}}$). The $\Delta_{p+ip}$ phase, on the other hand, restores $C_{3}$ symmetry (combined with a gauge transformation) so the energies are $C_{3}$-symmetric, and it exhibits a full gap ranging in size from $0.43\Delta$ along the $MK$ direction to $1.6\Delta$ along the $\Gamma M$ direction.
Optical conductivity–– Since we consider exclusively $k-$independent pairings, Peierls substitution couples the electromagnetic field $\vec{k}\to\vec{k}-e\vec{A}$ through the normal-state Hamiltonian $h_{k}$, and the conductivity can be computed using standard Kubo formalismAhn and Nagaosa (2021a):
$\sigma_{ij}=\frac{\pi e^{2}}{2\omega V}\sum_{m\neq n}v^{i}_{nm}v^{j}_{mn}f_{nm}\delta(\varepsilon_{n}-\varepsilon_{m}+\omega),$ (8)
where $e$ is the electron charge, $V$ is the volume of the system, $f_{nm}=f(\varepsilon_{n})-f(\varepsilon_{m})$ is the difference of Fermi functions, which depend on the energy $\varepsilon_{n}=E_{n}-\mu$ of band $n$, the chemical potential $\mu$ and the inverse temperature $\beta=1/k_{B}T$ measured in units of the Boltzmann constant $k_{B}$. We set $\beta=300\mathrm{eV}^{-1}$ ($T\simeq 3.87$K) [^1] It is important to note that the calculation of the optical conductivity has no information about the transition temperature of the superconductor and it does not require to be under the critical temperature. Therefore the temperature used here just plays the role of a small softening of the Fermi functions.. The velocity matrices $v^{i}_{mn}$ account for the transition amplitude between states $n,m$ and can be defined as:
$v^{i}_{nm}=\langle n|\frac{\partial H}{\partial k_{i}}|m\rangle=\langle n|\left(\begin{array}[]{cc}\partial_{k_{i}}h_{k}&0\\ 0&\partial_{k_{i}}\left[h^{T}_{-k}\right]\end{array}\right)|m\rangle.$ (11)
Note the optical conductivity of clean multiband superconductorsAhn and Nagaosa (2021a) as computed here can be reduced in size for some types of superconductors -but not suppressed- by inclusion of vertex correctionsWatanabe and Watanabe (2024), which are outside of the scope of this work.
Anisotropy in the nematic phase –– In the presence of $C_{3}$ symmetry, the optical conductivity satifies $\sigma_{xx}=\sigma_{yy}$. Since the two-component nematic pairing breaks this symmetry, we consider the difference $\sigma_{xx}-\sigma_{yy}$ as a probe of the nematic pairing. Fig. 3 shows this difference in the range of frequencies near the superconducting gap for both $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$. A sizable conductivity occurs due to transitions near the Fermi level, with significant differences between $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$ which can help determine the pairing anisotropy.
As a further optical signature of the pairing in multiorbital systems, we also consider higher frequency effects which occur due to transitions from the Fermi level BdG quasiparticles to unoccupied electron states, and from occupied electron states to BdG quasiparticles. These occur in the frequency range where the optical conductivity of the normal state is finite (see Fig. 1 (c)) and represent transfer of spectral weight due to superconductivity as discussed in Ref. [Ahn and Nagaosa, 2021b]. However, here we consider responses which are zero in the absence of pairing, due to $C_{3}$ and time-reversal symmetry. The observation of any optical conductivity is therefore a signature of multiorbital pairing, which is shown in the inset of Fig. 3. The anisotropy of the nematic pairing is reflected in this spectrum as well.
Figure 3: Anisotropy of the optical conductivity at low (main figures) and high (inset) frequencies for the nematic phases $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$ with $\Delta=0.1$ eV.
Hall conductivity in the chiral phase–– The chiral $E^{\prime}$ state given by the $\Delta_{p+ip}$ pairing breaks time-reversal symmetry, allowing a finite optical Hall conductivity $\sigma_{xy}^{H}=-\sigma_{yx}^{H}$. Fig. 4 shows a comparison of the optical Hall and regular conductivities in the chiral $E^{\prime}$ state with the pairing $\Delta_{p+ip}$ in the frequency range of the superconducting gap. The Hall conductivity is comparable to the interband optical responses $\sigma_{xx}=\sigma_{yy}$ in the material, but notably while both $\sigma_{xx}$ and $\sigma^{H}_{xy}$ exhibit a peak at around $1.6\Delta$ associated with the maximal gap size and a large JDOS of the electron and hole pockets along the $K\Gamma$ high-symmetry line (see Fig. 2(a)), the size of this peak is much smaller for $\sigma^{H}_{xy}$. At a second peak around $3.6\Delta$, however, exhibits a similar magnitude in $\sigma_{xx}$ and $\sigma^{H}_{xy}$ associated with the higher-frequency van Hove singularity along the $K\Gamma$ high-symmetry line (see Fig. 2 (c), dotted-dashed line). As explained in Appendix II.1, the anomalously small peak in $\sigma_{xy}^{H}$ occurs due to a cancellation of matrix elements from transitions originating from different bands.
Additionally, the $\Delta_{p+ip}$ also exhibits a finite contribution to the Hall conductivity at large frequencies due to the electronic transitions from Fermi surface BdG quasiparticles to higher-energy unpaired bands, shown in the inset of Fig. 4. This provides an additional high-energy fingerprint of the TRS breaking associated exclusively with the chiral phase.
Figure 4: Hall conductivity $\sigma^{H}_{xy}$ of the chiral superconducting state $\Delta_{p+ip}$ with $\Delta=0.1$ eV at low (main figure) and high (inset) frequencies. The diagonal component of the optical conductivity $\sigma_{xx}$ is shown for comparison at low frequencies.
Discussion–– In our work, we have predicted the optical conductivity of single layers of an H polytype of a TMD with $D_{3h}$ symmetry. These results most directly apply to single layers of NbSe2, where signatures of a nematic pairing channel have been observed in Refs. [Hamill et al., 2021] and [Cho et al., 2022]. Our results also bear relevance to 4Hb-TaS2, where superconductivity is dominated by the H layers in the T/H bulk structure. It should be noted this structure has an extra inversion center between the H layers and there are therefore two possibilities for inversion-even and odd versions of the $E^{\prime}$ state, which correspond to irreps $E_{2g}$ and $E_{1u}$ of the group $D_{6h}$ and should remain nearly degenerate. Finally, we speculate that our results may also be relevant to the superconducting state observed in CrBr3/NbSe2Kezilebieke et al. (2020), which coexists with ferromagnetism and shows chiral edge modes corresponding to a topological chiral superconductor.
[…]
In summary, we expect our work will help establish whether unconventional superconductivity in different TMD platforms is due to $E^{\prime}$ pairing, taking advantage of the multiorbital nature of their band structure with large Ising SOC which leads to unique optical conductivity signatures.
I Acknowledgements
F. J. is supported by Grant PID2021-128760NB0-I00 from the Spanish MCIN/AEI/10.13039/501100011033/FEDER, EU. M.-Á.S. was supported by EPSRC grant EP/X012239/1 and Programa Red Guipuzcoana de Ciencia Tecnología e Innovación No. 2021-CIEN-000070-01 Gipuzkoa Next.
II End Matter
II.1 Brillouin Zone distribution of the Hall response in $\Delta_{p+ip}$
In the chiral superconducting phase, the low-energy peak of the optical conductivity $\sigma_{xx}$ and the onset of the Hall conductivity is associated with the gap size, as discussed in the main text. The $\sigma_{xx}$ component exhibits a large peak around $1.6\Delta$ associated with the large JDOS between the closest bands to the Fermi level along the $K\Gamma$ high-symmetry lines and the positive-defined optical conductivity integrand over the Brillouin Zone (see Fig. 5 (a)). However, in the case of $\sigma^{H}_{xy}$ the velocity matrix elements connecting the closest bands to the Fermi level over the Brillouin Zone can be positive or negative (see Fig. 5 (b)). This leads to a cancellation of contributions that suppresses the peak that is present in the $\sigma_{xx}$ component due to the positive value of its velocity matrix elements over the whole Brillouin Zone.
The secondary peak at $\omega\sim 3.6\Delta$, associated with higher-frequency Van Hove singularity along $K\Gamma$ (see Fig. 2 (c), dashed-dotted line) is similar in size for $\sigma_{xx}$ and $\sigma^{H}_{xy}$ (see Fig. 4). This is due to the dominance of positive velocity matrix elements contributing to the Hall conductivity in that range of frequencies, much more similar to the $\sigma_{xx}$ case (Figs. 5 (c) and (d)) than in the primary peak case. The positive-integrand predominance in the secondary peak is due to the large positive contribution of the Van Hove singularity along $K\Gamma$, while having a limited presence of the negative-contributing bands (Fig. 5 (d)), which were more relevant in the primary peak energy range (Fig. 5 (b)).
Figure 5: Brillouin Zone-resolved optical conductivity $\sigma_{xx}$ (left column) and $\sigma^{H}_{xy}$ (right) column integrated between $\omega=1.3\Delta$ and $\omega=1.8\Delta$ (upper row) corresponding to the low-energy peak in the response in Fig. 4, and $\omega=3\Delta$ and $\omega=4\Delta$ (lower row) corresponding to the secondary peak in the response in Fig. 4.